OpenStudy (anonymous):

Simplify. 3 square root of 10 end root plus 7 square root of 15 end root minus 6 square root of 10 end root minus 4 square root of 15 3 square root of 10 end root minus 3 square root of 15 3 square root of 15 end root minus 3 square root of 10 3 square root of 30 end root minus 3 square root of 20 3 square root of 20 end root minus 3 square root of 30

OpenStudy (anonymous):

@lizz123

OpenStudy (lizz123):

http://gyazo.com/21f4004834f21304a8891ddafcc3700a

OpenStudy (mathstudent55):

Is this the problem? \(3\sqrt{10} + 7 \sqrt{15} - 6 \sqrt{10} - 4 \sqrt{15}\)

OpenStudy (anonymous):

???@mathstudent55

OpenStudy (mathstudent55):

Terms with sqrt(10) are like terms and can be combined together. Terms with sqrt(15) are like terms and can be combined together.

OpenStudy (anonymous):

ok @mathstudent55

OpenStudy (anonymous):

@lizz123 already told me that it was the 2nd one but thanks anyways

OpenStudy (anonymous):

she explained to me also :)

OpenStudy (mathstudent55):

\(\color{red}{3\sqrt{10}} + \color{green}{7 \sqrt{15}} + \color{red}{- 6 \sqrt{10}} + \color{green}{- 4 \sqrt{15}}\) \(= \color{red}{3\sqrt{10}} + \color{red}{- 6 \sqrt{10}} + \color{green}{7 \sqrt{15}} + \color{green}{- 4 \sqrt{15}}\) \(= \color{red}{3\sqrt{10}} \color{red}{- 6 \sqrt{10}} + \color{green}{7 \sqrt{15}} \color{green}{- 4 \sqrt{15}}\) \(= \color{red}{-3\sqrt{10}} + \color{green}{3 \sqrt{15}}\) \(= \color{green}{3 \sqrt{15}} \color{red}{-3\sqrt{10}} \) \(= 3 \sqrt{15} -3\sqrt{10} \)